### v1

# Grand-Canonical Symmetric Orbifold Theories

Abstract
In this letter, we study grand-canonical symmetric orbifolds of conformal field theories on the sphere. We show that a consistent operator product expansion can be defined, provided the Hilbert space is taken to be the direct sum of symmetric orbifolds of all degrees. This Hilbert space contains a tower of central operators of conformal dimension zero, one of which represents the central charge of the Virasoro algebra. Our construction thus provides a conformal field theory interpretation of the central charge operator $\mathcal{I}$ in $AdS_{3}$ string theory.
I Introduction
String theory on an $AdS_{3}\times S^{3}\times T^{4}$ background with one unit of NS-NS flux is dual to a grand-canonical ensemble of symmetric orbifold conformal field theories [1, 2, 3, 4, 5]. So far, these ensembles have been defined solely through their correlation functions. Up to a normalization factor, they read
$\langle\cdot\rangle_{p}:=\sum_{d}p^{d}\langle\cdot\rangle^{un}_{d}\,$, (1)
where $\langle\cdot\rangle^{un}_{d}$ denotes the unnormalized correlation function in the symmetric orbifold of degree $d$. The purpose of this letter is to investigate these ensembles further and show that they realize much of the structure of a single conformal field theory. In particular, we prove that operators in a grand-canonical symmetric orbifold admit an operator product expansion. Its consistency requires the Hilbert space to be enlarged beyond that of a single symmetric orbifold in the large $N$ limit. The Hilbert space transforms in a reducible representation of the Virasoro algebra that is not completely reducible. Its central charge is represented by a non-scalar operator, which we identify with the central charge operator $\mathcal{I}$ in the dual string theory. Our construction therefore sheds light on the nature of the operator $\mathcal{I}$ from a conformal field theory perspective.
I.1 Symmetric orbifolds
Symmetric orbifold theories are defined by tensoring $d$ copies of a two-dimensional seed conformal field theory and gauging the group $S_{d}$ that permutes the copies. The identification of configurations which differ only by a permutation of the copies introduces new sectors that obey twisted boundary conditions. In order to deal efficiently with these new sectors, we introduce twist fields $\sigma_{g}$ that generate the twisted vacua and thus modify the boundary conditions obeyed by the fields of the theory. We write
$X_{I}(e^{2\pi i}z)\sigma_{g}(0)=X_{g(I)}(z)\sigma_{g}(0)\,$, (2)
where $I$ labels the different copies of the field $X$. To obtain gauge-invariant twists, we sum over all permutations in a given conjugacy class $[g]$:
$\sigma_{[g]}(x):=\sum_{g\in[g]}\sigma_{g}(x)\,.$ (3)
The conformal dimension of a twist operator $\sigma_{k}$ associated to single cycles of length $k$ is given by [6]:
$h_{k}=\frac{c_{seed}}{24}\cdot\left(k-\frac{1}{k}\right)\,$, (4)
where $c_{seed}$ denotes the central charge of the seed theory. Furthermore, the Hilbert space of the symmetric orbifold theory is given by a direct sum over all sectors
$\mathcal{H}=\bigoplus_{[g]}\mathcal{H}_{[g]}\,.$ (5)
We note that the Hilbert space of a multi-cycle twisted sector is given by
$\mathcal{H}_{[g]}=\bigotimes_{n}(\underbrace{\mathcal{H}_{n}\otimes\dots\otimes\mathcal{H}_{n}}_{k_{n}\,\text{times}})^{S_{k_{n}}}\,$, (6)
where $k_{n}$ denotes the number of $n$ cycles in $g$. Moreover, the stress-energy tensor $T(z)$ of the orbifold theory is the sum of the stress tensors $T_{I}(z)$ of the individual copies. Accordingly, the central charge is $d\cdot c_{seed}$. Each sector $\mathcal{H}_{n}$ is generated by acting on the corresponding twisted ground state $\lvert\sigma_{n}\rangle$ with the fractional modes $(L_{m/n})_{m\in\mathbb{Z}}$ of the stress tensor in that sector [7, 8].
There are various ways to compute correlators of a two-dimensional symmetric orbifold theory, see [6, 9, 10, 8]. We briefly describe here the covering space approach developed in [6]. Given a manifold with non-trivial monodromies specified by a set of permutations $\{g_{1},\dots,g_{n}\}$, we construct the corresponding covering map $\Gamma$. We thereby trade $d$ fields obeying twisted boundary conditions for a unique single-valued field that lives on the covering surface. The correlators
$\langle\sigma_{g_{1}}(x_{1})\cdots\sigma_{g_{n}}(x_{n})\cdot O_{1}(y_{1})\cdots O_{m}(y_{m})\rangle_{d}:=\\ \frac{1}{d!}\int_{g_{i}-twisted}\mathcal{D}^{d}Xe^{-S[X]}O_{1}(y_{1})\cdots O_{m}(y_{m})\,.$ (7)
can be computed by lifting to the covering space, and gauge-invariant observables are then obtained by taking appropriate sums. This procedure requires a Weyl transformation to cancel the determinant of the covering map. The Weyl transformation modifies the measure of the path integral, conjuring a Liouville action which must be computed and appropriately regularized. The details of this procedure are not necessary to our discussion, and we therefore refer the reader to the original paper [6].
I.2 Road map
One motivation for studying grand-canonical ensembles of symmetric orbifold theories is the apparent lack of an operator product expansion. Consider indeed the following grand-canonical correlator:
$\langle\sigma_{k}(z)\sigma_{k}(0)\rangle_{p}:=\frac{1}{Z_{p}}\sum_{d\geq k}p^{d}\langle\sigma_{k}(z)\sigma_{k}(0)\rangle_{d}\,$, (8)
where $Z_{p}$ denotes the partition function of the grand-canonical theory. At each degree, one can take the operator product expansion of the twist operators. However, the corresponding OPE coefficients depend on the degree. For example, $C_{kk}^{\quad 0}(d)=\frac{d!}{k(d-k)!}$. Therefore, it is unclear whether an operator product expansion can be defined in the grand-canonical theory. To prove that such an expansion would be inconsistent, let us compute the grand-canonical two-point function explicitly:
$\langle\sigma_{k}(z)\sigma_{k}(0)\rangle_{p}=\frac{p^{k}Z_{seed}^{k}}{k}\frac{1}{z^{2h_{k}}}\,$, (9)
where $Z_{seed}$ denotes the partition function of the seed theory. We read off the OPE coefficient $g_{kk}^{\ \ 0}$:
$g_{kk}^{\ \ 0}=\frac{p^{k}Z_{seed}^{k}}{k}\,.$ (10)
This result is incompatible with certain four-point functions of the theory. Indeed, it implies that the contribution to a four-point function $\langle\sigma_{k}\sigma_{k}\sigma_{l}\sigma_{l}\rangle$ obtained from contracting the two $\sigma_{k}$ together and the two $\sigma_{l}$ together is
$\langle\sigma_{k}(x)\sigma_{k}(y)\sigma_{l}(z)\sigma_{l}(0)\rangle_{p}=\frac{p^{k+l}}{k\cdot l}\cdot Z_{seed}^{k+l}+\dots\,$, (11)
whereas a careful computation yields
$\frac{1}{k\cdot l}\frac{1}{Z_{p}}\sum_{d}p^{d}Z_{seed}^{d}\frac{d!}{(d-k)!(d-l)!}\,.$ (12)
This indicates that the grand-canonical theory lacks a straightforward operator product expansion.
Another line of inquiry that motivates the study of grand-canonical symmetric orbifolds comes from the non-scalar central charge in the algebra of currents of $AdS_{3}$ string theories with pure NS-NS flux [11].
The central charge operator $\mathcal{I}$ is the integrated vertex operator of winding number $w=1$, spin zero, and conformal dimension zero. It acts inside any fixed-genus correlator as a multiple of the identity [11], with a constant of proportionality set by the representations of the Virasoro algebra featured in the correlator [3]. The nature of its counterpart in the conformal field theory remains to be understood, and a general characterization likely requires further insight into the dual theory. That said, in settings where the dual field theory is known, the lack of an intrinsic conformal field theory definition of the central charge operator is striking. String theory on an $AdS_{3}\times S^{3}\times T^{4}$ background with one unit of NS-NS flux is dual to a grand-canonical ensemble of symmetric orbifolds with seed the torus superconformal theory. Where in this theory is the central charge operator found ?
An alternative way to describe the central charge operator $\mathcal{I}$ is as a differential operator with respect to the chemical potential [5]. As such, it generates an algebra of operators absent from the Hilbert space of the symmetric orbifold. Therefore, if the central charge operator of the string theory admits a field-theoretic dual, then its description requires the Hilbert space of the grand-canonical theory to be substantially enlarged. Following the intuition gained from the study of topological symmetric orbifolds [12, 13, 14], we will propose that the algebraic structure required for this enlargement is that of partial permutations [15]. A partial permutation is a permutation of a subset of $n$ elements. It is composed of a support $\mathcal{S}\subset\{1,\dots,n\}$ and a permutation $\rho$ of $\mathcal{S}$. We denote it as $(\mathcal{S},\rho)$, or alternatively with the Latin letter $r$. The product of two partial permutations is obtained by multiplying the permutations and concatenating their supports:
$(\mathcal{S}_{1},\rho_{1})\cdot(\mathcal{S}_{2},\rho_{2})=(\mathcal{S}_{1}\cup\mathcal{S}_{2},\rho_{1}\rho_{2})\,.$ (13)
Moreover, the symmetric group $S_{n}$ acts on partial permutations by conjugation. We write
$\pi\cdot(\mathcal{S},\rho)=(\pi\mathcal{S},\pi\rho\pi^{-1})\,.$ (14)
We define the orbit sum $A_{[r]}$ as the sum over all partial permutations in the conjugation orbit of $r$. These orbit sums are characterized by the size of their support and their cycle type. Our proposal is that in a grand-canonical theory, twist operators are indexed not by orbits of the infinite symmetric group, but rather by orbits of partial permutations. An immediate consequence of this proposal is that there are many central operators. In section III, we will identify these operators within the algebra generated by the dual of the central charge operator $\mathcal{I}$. In the next section, we focus on clarifying the structure of the grand-canonical theory: we build up the Hilbert space, explain how a consistent operator product expansion can be defined, and determine the representation of the Virasoro algebra.
II The Grand-Canonical Symmetric Orbifold Theory
II.1 The Hilbert Space
To construct the Fock space of quantum field theory, one must sum over all $n$-particle Hilbert spaces. Similarly, the Hilbert space of the grand-canonical symmetric orbifold theory is obtained by summing the Hilbert spaces corresponding to all degrees:
$\mathcal{H}=\bigoplus_{d\geq 0}\mathcal{H}_{d}\,.$ (15)
Each Hilbert space is itself the sum of twisted sectors. We may therefore write
$\mathcal{H}=\bigoplus_{[r]}\mathcal{H}_{[r]}\,$, (16)
where the sum runs over orbits of partial permutations of all sizes. We denote by $\alpha_{[r]}$ the twist operator that generates the ground state of the sector $\mathcal{H}_{[r]}$, and refer to it as a generalized twist operator. The partial permutation orbit of empty support $\alpha_{\emptyset}$ is the identity operator. It corresponds under the state-operator map to the vacuum of $\mathcal{H}_{0}$, which should therefore be regarded as the vacuum of the theory.
It is convenient to organize the Hilbert space in terms of the boundary conditions its fields obey. Given a partial permutation $r$ with minimal support $\mathcal{S}_{r}$, we define
$\hat{\mathcal{H}}_{r}:=\bigoplus_{k\geq 0}\mathcal{H}_{[(\mathcal{S}_{r}\cup 1^{k},r)]}\,.$ (17)
The Hilbert space $\mathcal{H}$ is the direct sum of the $\hat{\mathcal{H}}_{r}$.
II.2 Correlation Functions
We now explain how to compute disconnected correlation functions in the grand-canonical theory. We define a fixed-degree correlation function that involves a generalized twist field as
$\langle\alpha_{(\mathcal{S},\rho)}\cdots\rangle_{d}:=\langle\sigma_{\rho}\cdots\rangle^{un}_{d}\,.$ (18)
Although these correlators match, the corresponding gauge-invariant observables differ by a combinatorial factor. For example,
$\langle\alpha_{k,1}(z)\alpha_{k}(0)\rangle_{d}$ $=\frac{1}{k(d-k-1)!}\cdot\frac{Z_{seed}^{d}}{z^{2h_{k}}}\,$, (19)
which differs from $\langle\sigma_{k}\sigma_{k}\rangle_{d}$ by a factor dkd-k. This factor accounts for the larger number of elements entering the definition of $\alpha_{k,1}$, as compared to $\sigma_{k}$. In general, we find [15]:
$\langle\alpha_{[r]}(z)\cdot(\dots)\rangle_{d}=\binom{m_{1}(\rho)}{m_{1}(r)}\langle\sigma_{[\rho]}(z)\cdot(\dots)\rangle_{d}\,$, (20)
where $m_{1}(r)$ is the number of trivial elements in the support of the partial permutation $r$, and $m_{1}(\rho)$ the number of trivial elements in the corresponding permutation $\rho$ of $\{1,2,\dots,d\}$. Geometrically, the operator $\alpha_{[r]}$ selects $\lvert\mathcal{S}_{r}\rvert$ sheets out of the $d$ available in the covering space. Only those sheets contribute to connected correlators. This geometric picture is perhaps best understood from the dual string perspective, where trivial elements in the support of a partial permutation correspond to disjoint worldsheets in their vacuum state, see section III.4.
Additionally, we note that combinatorial differences in fixed-degree correlation functions can lead to drastically different behaviours once summed over all degrees. For example, the operator $\alpha_{1}$ behaves within any fixed-degree correlation function as a factor of the degree:
$\langle\alpha_{1}\cdots\rangle_{d}=d\cdot\langle\cdots\rangle_{d}\,.$ (21)
Nevertheless, its expectation value is non-trivial:
$\langle\alpha_{1}\rangle=pZ_{seed}\,.$ (22)
On the other hand, the generalized twist operators $\alpha_{k\geq 2}$ behave precisely as the standard $\sigma_{k}$ in all correlation functions, and we may therefore view the $\alpha_{k}$ as the standard twist operators in the grand-canonical theory.
Finally, we remark that the dependence of grand-canonical correlation functions on the positions of the fields always factors out of the sum over degrees, as exemplified by equation (9). It follows that the grand-canonical correlation functions inherit the conformal invariance of the fixed-degree correlators.
II.3 Operator Product Expansion
In order to take operator product expansions in the grand-canonical theory, we first compute at a given degree $d$, then replace each factor of $d$ by the operator $\alpha_{1}$. As we will see in section III.3, the product of $\alpha_{1}$ with any other twist operator can be calculated exactly. Multiplying out its powers leads to the desired expansion, which is manifestly independent of the degree. On the one hand, this immediately implies that the operator product expansion is consistent, because it factorizes out the grand-canonical sum. On the other hand, it implies the existence of a well-defined operator product expansion in the large $N$ limit of the fixed-degree symmetric orbifold. In the next section, we study the self-OPE of the stress-energy tensor of the theory. It provides a concrete example of the technique laid down above.
II.4 Virasoro Algebra and Central Charge
The stress-energy tensor of the grand-canonical theory coincides with the stress-energy tensor of a single symmetric orbifold in the large $N$ limit:
[…]
where $T_{I}$ denotes the stress-energy tensor of the $I$-th copy. Indeed, this is a conserved spin-$2$ tensor that generates conformal transformations, because the truncation to degree $d$ of $T(z)$ is the fixed-degree stress tensor $T_{d}$. We write
$\langle\left[\frac{1}{2\pi i}\int dz\,T(z)\epsilon(z),\Phi(w)\right]\cdots\rangle$ $=\sum_{d}p^{d}\langle\delta_{\epsilon}\Phi(w)\cdots\rangle_{d}$
$=\langle\delta_{\epsilon}\Phi(w)\cdots\rangle\,.$
The operator product expansion of the stress tensor with itself reads
$T(z)T(0)=\frac{(c_{seed}/2)\cdot\alpha_{1}}{(z-w)^{4}}+\frac{T(z)}{(z-w)^{2}}+\frac{\partial T(z)}{z-w}+\text{regular}\,.$ (24)
The central charge is represented non-trivially as the operator $c_{seed}\,\alpha_{1}$. The corresponding representation of the Virasoro algebra reads
$L_{n},L_{m}]=(n-m)L_{n+m}+\frac{c_{seed}\,\alpha_{1}}{12}(n^{3}-n)\delta_{n+m,0}\,$, (25)
where the $L_{n}$ are the Laurent modes of the stress-energy tensor, $L_{n}:=\int dz\,z^{n+1}T(z)$. An infinite-dimensional generalization of Schur’s lemma implies that the representation at hand is reducible, as it admits a non-scalar central element. In fact, the grand-canonical Hilbert space exhibits the structure of a descending filtration of reducible representations:
$\bigoplus_{d\geq 0}\mathcal{H}_{d}\supset\bigoplus_{d\geq 1}\mathcal{H}_{d}\supset\bigoplus_{d\geq 2}\mathcal{H}_{d}\supset\dots$ (26)
Nevertheless, the theory does not decompose into superselection sectors because the representation is not completely reducible. Indeed, the commutation relations of the Virasoro algebra generate the central charge operator, which is of degree one in the filtration. The Hilbert spaces $\hat{\mathcal{H}}_{r}$ defined in (17) are the indecomposable building blocks of the theory.
To conclude, we note that the appearance of a non-scalar central charge in the representation (25) is consistent with the findings of [16] for the spacetime Virasoro algebra of $AdS_{3}$. In the tensionless context, it leads to the identification of $\alpha_{1}$ as the counterpart of the central charge operator $\mathcal{I}$ of the dual string theory. The expectation value of $c_{seed}\,\alpha_{1}$ therefore computes the Brown-Henneaux central charge [17]. This allows us to precisely identify the relation between the string coupling and the chemical potential. The three-dimensional Newton constant for an $AdS_{3}\times S^{3}\times T^{4}$ background at $k=1$ reads [18]
$G_{N}^{(3)}=\frac{G_{N}^{(10)}}{\text{Vol}(S^{3})\cdot\text{Vol}(\mathbb{T}^{4})}=\frac{8\pi^{6}g_{s}^{2}R^{8}}{2\pi^{2}R^{3}\cdot(2\pi)^{4}v_{4}R^{4}}\,$, (27)
where $v_{4}=\frac{\text{Vol}(\mathbb{T}^{4})}{(2\pi R)^{4}}$. The Brown-Henneaux central charge is thus
[…]
which leads to
$pZ_{seed}=v_{4}\cdot g_{s}^{-2}\,$, (29)
consistent with [19]. We note however that the partition function of the seed conformal field theory has an ambiguity [20, 21].
II.5 A bona-fide conformal field theory ?
At this stage, let us summarize which conformal field theory axioms the grand-canonical symmetric orbifold satisfies. Firstly, the conformal invariance is manifest. Secondly, the state-operator correspondence and commutativity axioms hold by construction. Thirdly, we have shown that a consistent operator product expansion can be defined. Nevertheless, we do not expect grand-canonical ensembles to be well-defined on Riemann surfaces of genus $g\geq 2$, due to the rapid growth of the handle operator with the degree. In fact, these theories fail to satisfy the sewing axiom of [22]. Moreover, their spectrum features reducible but indecomposable representations. Such representations are emblematic of logarithmic conformal field theories. In these theories, the dilation operator is not diagonalizable, which leads to logarithmic singularities for the correlation functions [23]. The central charge, on the other hand, is represented by a multiple of the identity. Interestingly, the grand-canonical symmetric orbifold satisfies the same axioms as a logarithmic conformal field theory on the sphere, but shares neither of the properties described above. The takeaway is that the axiom of decomposability can be violated in qualitatively distinct ways.
III Central Operators
The generalized twist operator $\alpha_{1}$ is of special interest. Indeed, it represents the central charge of the Virasoro algebra of the grand-canonical theory. Furthermore, together with the standard twist operators $\alpha_{k}$ and the seed excitations, it generates the entire spectrum of the theory. In this section, we study some of its properties.
III.1 The Central Charge Operator
The operator $\alpha_{1}(x)$ is independent of its position $x$. This follows immediately from equation (20). We will therefore consistently omit its insertion point. Furthermore, $\alpha_{1}$ acts on unnormalized correlators as the differential operator $\partial_{\log p}$:
$\langle\alpha_{1}\cdot(\dots)\rangle^{un}$ $:=\sum_{d\geq 1}d\,p^{d}\langle(\dots)\rangle^{un}_{d}$ (30)
$=p\partial_{p}\langle(\dots)\rangle^{un}\,.$ (31)
This behaviour is consistent with that of the central charge operator $\mathcal{I}$ of an $AdS_{3}$ string [21, 19]. In particular, it implies that inserting the latter operator in a disconnected, fixed-degree correlator multiplies it by a factor of the degree.
III.2 The Universal Subalgebra
Let us now study the tower of operators $\alpha_{1^{k}}$, all of which are once again independent of their insertion point. Moreover, they are central, meaning that they commute with every operator in the theory. Their vacuum expectation value is
$\langle\alpha_{1^{k}}\rangle$ $=\frac{p^{k}Z_{seed}^{k}}{k!}\,.$ (32)
Furthermore, their action within any correlation function reads
$\langle\alpha_{1^{k}}\cdot(\dots)\rangle^{un}$ $:=\sum_{d\geq 1}\binom{d}{k}p^{d}\langle(\dots)\rangle^{un}_{d}\,$, (33)
$=\frac{p^{k}\partial_{p}^{k}}{k!}\langle(\dots)\rangle^{un}\,.$ (34)
Therefore, we identify
$\alpha_{1^{k}}\equiv\frac{p^{k}\partial_{p}^{k}}{k!}\,.$ (35)
The algebra of central operators is isomorphic to the algebra of differential operators $p^{k}\partial_{p}^{k}/k!$. It can be obtained by induction. To begin with, we compute
$\alpha_{1^{k+1}}\alpha_{1^{l}}=\frac{1}{k+1}\,(\alpha_{1}-k)\cdot\alpha_{1^{k}}\alpha_{1^{l}}\,.$ (36)
We denote by $a_{kl}^{\ \ m}$ the coefficients of $\alpha_{1^{m}}$ in the product $\alpha_{1^{k}}\cdot\alpha_{1^{l}}$. Equation (36) leads to the following recursion relation for the coefficients $a_{kl}^{\ \ m}$:
$a_{k+1,l}^{\quad\quad n}=\frac{1}{k+1}\left[(n-k)a_{kl}^{\ \ n}+na_{kl}^{\ \ n-1}\right]\,$, (37)
By solving this relation, we obtain the following algebra of central operators
$\alpha_{1^{k}}\alpha_{1^{l}}=\sum_{m=0}^{\min(k,l)}\frac{(k+l-m)!}{m!(k-m)!(l-m)!}\;\alpha_{1^{k+l-m}}\,.$ (38)
By construction, this algebra is isomorphic to the freely generated algebra of the single element $\mathcal{I}$. The content of equation (38) is combinatorial: in the basis most appropriate to our problem, the structure constants take the form above. Furthermore, the fact that operators generated by $\mathcal{I}$ define an algebra is consistent with the independence of this operator on its position in spacetime.
[…]
The formula (39) effectively identifies the polynomial $P_{n}(x)$ that appeared in [3].
Finally, we propose to view the existence of this algebra as a diagnostic for determining whether a theory is grand-canonical: a conformal field theory is grand-canonical if and only if it contains an algebra freely generated by a single central operator.
III.3 OPE with the Central Charge Operator
The action of the central charge operator within any fixed-degree correlator is purely combinatorial. Therefore, the operator product expansion of any power of the central charge with a generalized twist reduces to their product as orbits of partial permutations [15]. To compute it, we first write
$\alpha_{k}\cdot\alpha_{1^{l}}=\sum_{i=0}^{\min(k,l)}\binom{k}{i}\alpha_{k,1^{l-i}}\,.$ (40)
We then expand $\alpha_{1}^{n}$ using equation (31), plug the above expression in the product and regroup terms. For simplicity, we will assume that $k\geq n$. We find
[…]
Stirling numbers of the second kind are generated by the exponential function $\frac{1}{l!}(e^{x}-1)^{l}$. Therefore, the coefficient $a^{(k,n)}_{p}$ of $\alpha_{k,1^{p}}$ is given by
$a^{(k,n)}_{p}$ $=n![x^{n}]e^{kx}(e^{x}-1)^{p}$ (42)
$=\sum_{j=0}^{p}(-1)^{p-j}\binom{p}{j}(k+j)^{n}\,.$ (43)
The simplest example is the operator product expansion of the central charge operator with a single cycle twist. It reads
$\alpha_{1}\alpha_{k}=\alpha_{k,1}+k\alpha_{k}\,$, (44)
with $k\geq 2$. A geometric interpretation of equation (44) is that $\alpha_{1}$ either marks an existing sheet or creates a new one. Holographically, the latter corresponds to adding a disjoint worldsheet in its vacuum state. Indeed, the Hilbert space (16) distinguishes between vacuum configurations with different numbers of worldsheets. A key takeaway of our construction is that for the boundary description to exhibit a field theory structure, it must holographically encode a multi-string Hilbert space.
III.4 Connected Correlation Functions
Connected correlation functions are obtained by restricting the sum over covering surfaces to connected coverings. The surfaces relevant to this definition are determined solely by the partial permutations that appear in the correlator. Hence, we say that a covering space is connected if the corresponding set of partial permutations acts transitively on its support. For example, $\langle\alpha_{(12)(3)}\alpha_{(23)}\rangle$ gives rise to a connected covering, whereas $\langle\alpha_{(12)(4)}\alpha_{(23)}\rangle$ does not. Consequently, the action of the central charge operator in connected correlators reads
$\langle\alpha_{1}\cdot\prod_{i}\alpha_{r_{i}}\rangle_{d}^{conn}=N_{c}\cdot\langle\prod_{i}\alpha_{r_{i}}\rangle_{d}^{conn}\,$, (45)
where $N_{c}$ is the number of connected sheets in the surface defined by the partial permutations $r_{i}$. Indeed, as previously explained, the operator $\alpha_{1}$ either marks a sheet or adds a new one, and only the former action contributes in the connected case. In fact, this behaviour follows directly from our identification of $\alpha_{1}$ as the dual of the central charge operator $\mathcal{I}$, as shown in [11]. It is also consistent with our interpretation of $\alpha_{1}$ as the analogue of a particle number.
IV Outlook
In this letter, we have shown that grand-canonical ensembles of symmetric orbifold conformal field theories can be formulated in such a way that they satisfy the same axioms as a logarithmic conformal field theory on the sphere. The appropriate formulation is best suited to describe a second-quantized Hilbert space for the string theory. In this context, we have provided a conformal field theory interpretation of the central charge operator of an $AdS_{3}$ string theory at the tensionless point. It would be interesting to understand the field-theoretic origin of the central charge operator beyond the unit-flux case. String theory on an $AdS_{3}\times X$ background with pure NS-NS flux is believed to be dual to a deformed symmetric orbifold in the grand-canonical ensemble [25, 26]. We therefore expect the ideas developed here to be of relevance in the more general case.
Another avenue for future work would be to generalize our construction to symmetric orbifolds with boundaries. This may shed light on some grand-canonical aspects of holographic interfaces in $AdS_{3}/CFT_{2}$ [27, 28, 29, 30]. A first step in this direction was taken in [31], which explores grand-canonical Hurwitz theories with boundaries.
Acknowledgements.
It is a pleasure to thank Jan Troost for many stimulating discussions and for comments on an early version of the manuscript. I am also grateful to António Antunes and Eric Perlmutter for insightful questions.